A thin stream of water flows smoothly from a fauce
Answer #1

A1*V1 = A2*V2

V2 = A1*V1 /A2

A1 / A2 = (1/ 0.751)2

V2 =(1/ 0.751)2 1.55 = 2.74822

v^2 =u^2 +2*g*h

h= (v^2 -u^2 ) / 2*9.8 = 26.27 cm